# LADDER: Self-Improving LLMs Through Recursive Problem Decomposition

### v1

@@ Abstract
We introduce LADDER (Learning through Autonomous Difficulty-Driven Example Recursion), a framework enabling LLMs to autonomously improve their problem-solving capabilities through self-guided learning. By recursively generating and solving progressively simpler variants of complex problems, LADDER enables models to progressively learn through reinforcement learning how to solve harder problems. This self-improvement process is guided by verifiable reward signals, allowing the model to assess its solutions. Unlike prior approaches requiring curated datasets or human feedback, LADDER leverages the model’s own capabilities to easier variants of sample questions. We demonstrate LADDER’s effectiveness on mathematical integration tasks, where it improves a Llama 3B model’s accuracy from 1% to 82% on undergraduate-level problems and enables a 7B parameter model to achieve state-of-the-art performance - 73% - on the MIT Integration Bee examination for it’s model size. We also introduce TTRL (Test-Time Reinforcement Learning), a method that generates variants of test problems at inference time and applies reinforcement learning to further improve performance. By further creating and solving related problems during testing, TTRL enables the 7B model to achieve a score of 90%, surpassing o1. These results showcase how strategic self-directed learning can achieve significant capability improvements without relying on architectural scaling or human supervision.
@@ 1 Introduction
Building on LADDER’s self-improvement framework, we also propose Test-Time Reinforcement Learning (TTRL), a novel approach that extends these principles to inference time. TTRL dynamically generates problem variants during test-time and applies reinforcement learning to refine the model’s solutions, effectively creating a micro-learning process for each test instance. By leveraging the same verification mechanisms used in training, TTRL enables the model to further improve its performance. When applied to the 2025 MIT Integration Bee, TTRL boosts accuracy from 73% - with just LADDER - to 90%, demonstrating how scaling test-time compute through strategic problem decomposition can yield substantial performance improvements. We achieve state of the art accuracy, outperforming significantly larger models, such as OpenAI’s o1.
@@ 3.1.4 Reinforcement Learning Protocol
We decided to employ Group Relative Policy Optimization (GRPO) in both LADDER and TTRL for two reasons [11]. First, GRPO does not use a seperate critic model and instead estimates the baseline from group scores, improving efficiency and reducing memory. For each question $q$, GRPO samples a group of outputs ${o_{1},o_{2},...,o_{G}}$ from the old policy $\pi_{\theta_{old}}$ and then optimizes the policy model $\pi_{\theta}$ by maximizing the following objective:
@@ 4.3 MIT Integration Bee (LADDER + TTRL)
After applying LADDER to the 7B base model, we further employed Test-Time Reinforcement Learning (TTRL) for 100 steps on questions that LADDER failed to answer correctly. Of the remaining incorrect responses, TTRL successfully solved 3 additional problems, increasing performance from 73% to 90% (Figure 6). This improvement places our model’s performance well above the typical qualification threshold of 73% for the MIT Integration Bee, setting a new state-of-the-art for LLMs on this benchmark.
@@ 6 Conclusion
Looking ahead, this research opens new paths for developing more autonomous and self-improving AI systems. The principles demonstrated here could extend to any domain with clear verification mechanisms, from program synthesis to formal theorem proving. By providing a framework for models to bootstrap their own capabilities through carefully constructed learning trajectories, we move closer to AI systems that can systematically extend their abilities into increasingly complex domains.

### v2

@@ Abstract
We introduce LADDER (Learning through Autonomous Difficulty-Driven Example Recursion), a framework which enables Large Language Models to autonomously improve their problem-solving capabilities through self-guided learning by recursively generating and solving progressively simpler variants of complex problems. Unlike prior approaches that require curated datasets or human feedback, LADDER leverages a model’s own capabilities to generate easier question variants. We demonstrate LADDER’s effectiveness on the subject of mathematical integration, improving Llama 3.2 3B’s accuracy from 1% to 82% on undergraduate-level problems and enabling Qwen2.5 7B Deepseek-R1 Distilled to achieve 73% on the MIT Integration Bee qualifying examination. We also introduce TTRL (Test-Time Reinforcement Learning), where we perform reinforcement learning on variants of test problems at inference time. TTRL enables Qwen2.5 7B Deepseek-R1 Distilled to achieve a state-of-the-art score of 90% on the MIT Integration Bee qualifying examination, surpassing OpenAI o1’s performance. These results show how self-directed strategic learning can achieve significant capability improvements without relying on architectural scaling or human supervision.
@@ 1 Introduction
Building on LADDER’s self-improvement framework, we propose Test-Time Reinforcement Learning (TTRL), a novel approach that extends these principles to inference time. TTRL dynamically generates problem variants during test-time and applies reinforcement learning to refine the model’s solutions, effectively creating a micro-learning process for each test instance. By leveraging the same verification mechanisms used in training, TTRL enables the model to further improve its performance. When applied to the 2025 MIT Integration Bee, TTRL boosts accuracy from 73% - with just LADDER - to 90%, demonstrating how scaling test-time compute through strategic problem decomposition can yield substantial performance improvements. We achieve state of the art accuracy, outperforming significantly larger models, such as OpenAI’s o1.
@@ 3.1.4 Reinforcement Learning Protocol
We decided to employ Group Relative Policy Optimization (GRPO) in both LADDER and TTRL [11]. GRPO does not use a separate critic model and instead estimates the baseline from group scores, improving efficiency and reducing memory. For each question $q$, GRPO samples a group of outputs ${o_{1},o_{2},...,o_{G}}$ from the old policy $\pi_{\theta_{old}}$ and then optimizes the policy model $\pi_{\theta}$ by maximizing the following objective:
@@ 4.3 MIT Integration Bee (LADDER + TTRL)
After applying LADDER to the 7B base model, we further employed Test-Time Reinforcement Learning (TTRL) for 100 steps on questions that LADDER failed to answer correctly. Of the remaining incorrect responses, TTRL successfully solved 3 additional problems, increasing performance from 73% to 90% (Figure 6). This improvement places our model’s performance well above the typical qualification threshold of 73% for the MIT Integration Bee, setting a new state-of-the-art for mid-sized LLMs on this benchmark.
@@ 6 Conclusion
The principles demonstrated in this paper could likely extend to any domain with clear verification mechanisms, from program synthesis to formal theorem proving. By providing a framework for models to bootstrap their own capabilities through carefully constructed learning trajectories, we move closer to AI systems that can systematically extend their abilities into increasingly complex domains.